OpenStudy (anonymous):

Use the distributive property to factor the expression below. -15x^2y^2 + 25xy^2

OpenStudy (paxpolaris):

Distributive property: \[a(P+Q) = aP + aQ\]

OpenStudy (anonymous):

huh D;?

OpenStudy (anonymous):

-15x^2y^2 + 25xy^2 we find factors of each term -15x^2y^2 = -3*5*x*x*y*y and 25xy^2 = 5*5*x*y*y see what is common (GCF) it is 5*x*y*y = 5xy^2 so we get factor as 5xy^2 (-3x + 5)